OpenStudy (anonymous):

find the derivative sin(5x^2+2x+1)^3 please help

OpenStudy (anonymous):

You would need to use the power rule and the chain rule.

OpenStudy (anonymous):

Oh and the derivative of sin

OpenStudy (anonymous):

So...\[dy = d(\sin(5x^2+2x+1)^3)\] First it's the power rule and chain rule:\[dy = 3\sin(5x^2+2x+1)^2 d(\sin(5x^2+2x+1))\]Then the derivative of sin and the chain rule:\[dy = 3\sin(5x^2+2x+1)^2 \cos(5x^2+2x+1)d(5x^2+2x+1)\]Then just the power rule:\[dy = 3\sin(5x^2+2x+1)^2 \cos(5x^2+2x+1)(10x+2)dx\]Then simplify

OpenStudy (anonymous):

thank you so much!